I NEED HELP! This problem is not coming out right. Here it goes...
Given the function k describe by k(x)=10x find each of the following:
(a) k(-17)
(b) k(8)
(c) k(34)
the answer k(-17) = ?????
PLEASE.....PLEASE....

Answer by rapaljer(4671)

If k(x)=10x, then you substitute the value inside the parentheses in place of x.
So, if
k(x)=10(x)
k(-17) = 10(-17)=-170

Likewise,
k(8)=10(8)=80
k(34)=10(34)=340
etc.

I have a complete explanation for this topic on my website. Click on my tutor name "rapaljer" anywhere in algebra.com. Look for the second link on my homepage which is "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", then "Chapter 4 Graphing." This lesson is "Section 4.09 Functional Notation."

R^2